# Synergistic Bleeding Region and Point Detection in Laparoscopic Surgical Videos

### v2

@@ 4.4 Loss Function
The total loss function of BlooDet consists of mask loss $\mathcal{L}_{mask}$ and edge loss $\mathcal{L}_{edge}$ from mask branch, as well as point loss $\mathcal{L}_{point}$ and score loss $\mathcal{L}_{score}$ from point branch:
$\mathcal{L}=\lambda_{m}\mathcal{L}_{mask}+\lambda_{e}\mathcal{L}_{edge}+\lambda_{s}\mathcal{L}_{score}+\lambda_{p}\mathcal{L}_{point}.$ (9)
@@
Both $\mathcal{L}_{mask}$ and $\mathcal{L}_{edge}$ are computed as a combination of Focal loss [20] and Dice loss [25]. In addition, $\mathcal{L}_{point}$ employs the smooth L1 loss for point-level supervision:
$\mathcal{L}_{point}=\sum_{i=1}^{N}\mathbf{1}_{\{p_{i}\neq[0,0]\}}\cdot L1(\hat{y}_{\text{p}}^{i},y_{\text{p}}^{i}),$ (10)
where $\hat{y}_{\text{p}}^{i}$ is the predicted point location and $y_{\text{p}}^{i}$ is the ground-truth point location. $\mathbf{1}_{\{p_{i}\neq[0,0]\}}$ denotes an indicator function that ensures the loss is calculated only when the point is not zero. Besides, $\mathcal{L}_{score}$ is the binary cross-entropy loss for point existence in point branch. $\lambda_{m}$, $\lambda_{e}$, $\lambda_{s}$, and $\lambda_{p}$ are empirically set to 1, 1, 1, and 0.5, respectively, to balance the total loss function.
@@ 5.1 Datasets and Evaluation Metrics
Datasets. Since the tasks of the laparoscopic bleeding region and bleeding point cooperative detection are proposed for the first time, we adopt the proposed SurgBlood dataset to train and test our method and related comparative methods. We randomly split a total of 95 video clips into two sets: 75 for training and the remaining 20 for testing.
Table 1: Overall comparison with well-known task-related methods on SurgBlood test set. SAM 2 denotes SAM 2_base with an additional point head.
Types Methods Volumes Bleeding Region Metrics Bleeding Point Metrics Params (M) $\downarrow$
IoU $\uparrow$ Dice $\uparrow$ PCK-2% $\uparrow$ PCK-5% $\uparrow$ PCK-10% $\uparrow$
Region-level Swin-UNet [4] ECCV’22 41.31 58.47 - - - 27.2
SAM [16] ICCV’23 40.43 57.49 - - - 93.7
SAM-Adapter [6] ICCV’23 54.80 70.80 - - - 93.8
MemSAM [10] CVPR’24 55.34 71.28 - - - 133.3
SAM 2 [29] ICLR’25 63.51 77.68 - - - 80.8
SAM2-Adapter [7] arXiv’24 64.23 77.95 - - - 88.8
Point-level HRNet [32] CVPR’19 - - 3.13 15.98 44.31 63.6
SimCC [19] ECCV’22 - - 2.14 14.99 46.95 66.3
GTPT [39] ECCV’24 - - 2.80 13.01 38.38 16.7
D-CeLR [8] ECCV’24 - - 5.10 27.67 60.13 53.4
Multi-task PAINet [9] MICCAI’23 44.14 61.24 2.47 15.48 48.43 13.6
PitSurgRT [23] IJCARS’24 30.48 46.72 2.47 13.84 41.68 67.3
SAM 2 [29] ICLR’25 50.93 67.49 12.35 41.68 71.99 81.0
BlooDet (Ours) - 64.88 78.70 18.62 55.85 83.69 91.6
Evaluation Metrics. Following previous studies [29, 9], we adopt the Intersection over Union (IoU) and Dice Coefficient (Dice) metrics to evaluate bleeding region detection performance. For bleeding points, the Percentage of Correct Keypoints (PCK) metric is used to measure localization accuracy. Unlike the 10% to 40% threshold range applied in [9, 23] for anatomical structure centroids, we adopt a narrower threshold range of 2%–10% to ensure greater assessment, i.e., PCK-2%, PCK-5%, and PCK-10%. This is due to the requirement for higher precision and lower tolerance of bleeding point detection in laparoscopic surgery.
@@ 5.2 Implementation Details
Our framework is implemented on a single RTX 4090 GPU. During training, we input eight consecutive frames in an online manner with resolution resized to 512$\times$512 pixels. The image encoder of BlooDet is initialized with pre-trained weights from SAM 2_base [29]. Additionally, we utilize a frozen PWC-Net [31] to compute inter-frame optical flow. No data augmentation is applied during data loading. The maximum learning rate for the image encoder is set to 5e-6, while other parts are trained with 5e-4. We employ the Adam optimizer with a warm-up strategy and linear decay, training for 20 epochs. During inference, we perform frame-by-frame inference in line with SAM 2.
@@
5.3 Performance on SurgBlood Benchmark
We evaluate 12 task-related methods for bleeding region and point detection to build a comprehensive benchmark on SurgBlood, including multi-task detection models [9, 23], video- and image-level object segmentation methods [4, 16, 29, 6, 10, 7], and pose-based point detection methods [32, 19, 39, 8]. For fairness, all methods use the official code and adapt only the head to fit bleeding tasks.
Quantitative Evaluation. Table 1 presents the performance of our framework and other comparison methods on SurgBlood test set for bleeding region and point detection. Our framework outperforms competitors across both tasks, even surpassing models designed for a single task. Benefiting from edge-enhanced prompts and bleeding point guidance, BlooDet achieves notable improvements in bleeding region detection. For bleeding point detection, BlooDet significantly surpasses other methods, with 6.27% and 14.17% improvement in PCK-2% and PCK-5% metrics, respectively. Besides, our framework obtains superior performance while adding a few parameters.
Qualitative Evaluation. Fig. 5 shows a visual comparison of our approach with other multi-task frameworks. We can see that BlooDet provides greater stability and consistency in detecting bleeding regions and points. In complex surgical environments with low contrast, competitors tend to be disturbed by surrounding noise. In contrast, our method ensures robust detection across consecutive frames via spatial-temporal modeling and co-guidance.
@@ 5.3 Performance on SurgBlood Benchmark
Table 2: Ablations for key components of BlooDet. EG and PMM denote edge generator and point memory modeling modules.
@@
Figure 6: Influence of edge generator for bleeding region detection.
@@
Global 15.49 49.59 82.00
Figure 7: Ablations for optical flow operation.
@@
Figure 9: Effect of mutual guidance between mask and point branches of BlooDet.
5.4 Ablation Analysis
Contributions of Key Component. Table 2 illustrates the contribution of key components in BlooDet. The experimental results show that point memory modeling contributes significantly to detecting bleeding points, e.g., improving the PCK-2% score by about 10.4%. Moreover, our edge generator provides effective edge prompt embedding, which greatly improves the accuracy of bleeding region detection. In short, each key component contributes positively to model performance.
Ablations for Edge Generator. We investigate the effect of different edge generator designs embedded in the mask branch. As exhibited in Table 9, the Wavelet Laplacian filter shows a strong response to edge clues, effectively mitigating interference from complex backgrounds in laparoscopic scenes. Meanwhile, integrating high-resolution features further enhances the quality of edge maps.
Optical Flow Operation Design. In point memory modeling, we adopt reversed mask maps in conjunction with optical flow maps to estimate the average camera displacement. To validate the effect of this design, we ablate the impact of focusing on different regions in mask maps for bleeding point detection. Table 9 indicates that using the foreground region leads to inferior performance. It may be explained by the poor stability of optical flow in the rapidly changing bleeding area. In contrast, utilizing the background enables more stable motion modeling and point localization.
Effect of Bidirectional Guidance. To validate the effectiveness of our cross-branch bidirectional guidance, Fig. 9 ablates the impact of point prompt from point branch as well as mask map and mask memory from mask branch. The results indicate that the point prompt contributes to bleeding region detection. For the point branch, the mask maps from previous frames assist with the optical flow operation, while mask memory fosters point memory modeling to improve the accuracy of point localization. It also illustrates that the proposed dual-branch collaborative structure is superior to a single-branch design. The attention maps from the mask and point decoders visualized in Fig. 9 also verify the effect of bidirectional guidance.
@@ 6 Conclusion
This work advances the intelligent detection of bleeding regions and bleeding points in laparoscopic surgical videos. We first contribute an open-source dataset for bleeding detection in real surgical scenarios, SurgBlood, to facilitate benchmark construction. Accordingly, we design a dual-task synergistic online framework called BlooDet, which assembles mask and point branches in a bidirectional guidance structure, and exploits an edge generator and point memory modeling to enhance the adaptive prompting scheme. Extensive experimental results demonstrate that our method outperforms existing models for both bleeding tasks. We believe this study can facilitate research in intelligent surgical assistance, reducing intraoperative decision-making risks and improving clinical outcomes.

### v3

@@ 4.4 Objective Function
The loss $\mathcal{L}_{\mathtt{m}}$ of mask branch consists of the region loss $\mathcal{L}_{\mathtt{r}}$ and edge loss $\mathcal{L}_{\mathtt{e}}$:
$\mathcal{L}_{\mathtt{m}}=\lambda_{\mathtt{r}}\mathcal{L}_{\mathtt{r}}+\lambda_{\mathtt{e}}\mathcal{L}_{\mathtt{e}}.$ (12)
@@
$\mathcal{L}_{\mathtt{r}}$ and $\mathcal{L}_{\mathtt{e}}$ are computed as a combination of Focal loss [lin2017focal] and Dice loss [milletari2016v]. In addition, the point branch loss $\mathcal{L}_{\mathtt{p}}$ consists of the point loss $\mathcal{L}_{\mathcal{P}}$ and the scoring loss $\mathcal{L}_{\mathtt{s}}$:
$\mathcal{L}_{\mathtt{p}}=\lambda_{\mathcal{P}}\mathcal{L}_{\mathcal{P}}+\lambda_{\mathtt{s}}\mathcal{L}_{\mathtt{s}}.$ (13)
$\mathcal{L}_{\mathcal{P}}$ employs the smooth L1 loss for point-level supervision. $\mathcal{L}_{\mathtt{s}}$ is the binary cross-entropy loss for point existence. $\lambda_{\mathtt{r}}$, $\lambda_{\mathtt{e}}$, $\lambda_{\mathtt{s}}$, and $\lambda_{\mathcal{P}}$ are set to 1, 1, 1, and 0.5. The training process of BlooDet can refer to Algorithm. 1 in Supp. E.
@@ 5.1 Datasets and Evaluation Metrics
Datasets. Since the tasks of the laparoscopic bleeding region and bleeding point cooperative detection are proposed for the first time, we adopt the SurgBlood dataset to train and test our method and related comparative methods. We randomly split a total of 95 video clips into two sets: 75 for training and the remaining 20 for testing. For further evaluation, we assess the region-level detection performance on HemoSet [miao2024hemoset], and provide additional comparative results on Rabbani et al. [rabbani2022video] and SUN-SEG [ji2022video] in Supp. F.
Evaluation Metrics. Following previous studies [ravi2024sam, das2023multi], we adopt the Intersection over Union (IoU) and Dice Coefficient (Dice) metrics to evaluate bleeding region detection performance. For bleeding points, the Percentage of Correct Keypoints (PCK) metric is used to measure localization accuracy. Unlike the 10% to 40% threshold range applied in [das2023multi, mao2024pitsurgrt] for anatomical structure centroids, we adopt a narrower threshold range of 2%–10% to ensure greater assessment, i.e., PCK-2%, PCK-5%, and PCK-10%. This is due to the requirement for higher precision and lower tolerance of bleeding point detection in laparoscopic surgery.
Table 3: Comparison with region-specific methods on SurgBlood.
Methods Volumes IoU $\uparrow$ Dice $\uparrow$
Swin-UNet [swinunet] ECCV’22 $41.31\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $58.47\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$
SAM [kirillov2023segment] ICCV’23 $40.43\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.13}}$ $57.49\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.15}}$
MemSAM [deng2024memsam] CVPR’24 $55.34\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.23}}$ $71.28\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.13}}$
STDDNet [chen2025stddnet] ICCV’25 $50.42\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.17}}$ $67.04\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.15}}$
SAM 2 [ravi2024sam] ICLR’25 $63.51\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.11}}$ $77.68\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.19}}$
[…]
@@ 5.2 Implementation Details
Our framework is implemented on a single RTX 4090 GPU. During training, we input eight consecutive frames in an online manner with resolution resized to $\times512$ pixels. The image encoder of BlooDet is initialized with pre-trained weights from SAM 2_base [ravi2024sam]. Additionally, we utilize a frozen PWC-Net [sun2018pwc] to compute inter-frame optical flow. No data augmentation is applied during data loading. The maximum learning rate for the image encoder is set to 5e-6, while other parts are trained with 5e-4. We employ the Adam optimizer with a warm-up strategy and linear decay, training for 20 epochs. During inference, we perform frame-by-frame inference in line with SAM 2.
5.3 Performance Comparison
We evaluate 13 task-related methods for bleeding region and point detection to build a comprehensive benchmark on SurgBlood, including multi-task detection models [das2023multi, mao2024pitsurgrt, mao2025consistnet], region-level object segmentation methods [swinunet, kirillov2023segment, ravi2024sam, deng2024memsam, chen2025stddnet], and point-level detection methods [sun2019deep, li2022simcc, wang2024gtpt, dai2024cephalometric, feng2025uncertainty]. Besides, we conduct zero-shot testing on the HemoSet [miao2024hemoset] dataset using pre-trained models from SurgBlood.
@@ 5.3 Performance Comparison
Quantitative Evaluation. We display the performance of our framework and comparison methods for bleeding region and point detection in Table 2. Region-level and point-level models respectively append point/region prediction heads for dual-task learning. We can see that our framework outperforms competitors across both tasks on SurgBlood and HemoSet [miao2024hemoset] datasets. Benefiting from edge-enhanced prompts and bleeding point guidance, BlooDet achieves notable improvements in bleeding region detection. For bleeding point localization, BlooDet significantly surpasses other methods thanks to mutual coupling guidance. Moreover, BlooDet obtains superior accuracy with proper parameters.
We also compare with task-specific models on individual tasks. As shown in Table 3 and Table 4, BlooDet maintains state-of-the-art performance in both bleeding region detection and point localization. Comparing the results in Table 2 further reveals that task-specific models impact the performance of the original task during dual-task learning, illustrating the effectiveness of our cross-branch guidance and alternating optimization strategy in multi-task learning. More fine-grained comparisons can be found in Supp. F.
Table 4: Comparison with point-specific methods on SurgBlood.
Methods Volumes PCK-2% $\uparrow$ PCK-5% $\uparrow$ PCK-10% $\uparrow$
[…]
SimCC [li2022simcc] ECCV’22 $2.14\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.18}}$ $14.99\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.22}}$ $46.95\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.11}}$
GTPT [wang2024gtpt] ECCV’24 $2.80\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.15}}$ $13.01\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.20}}$ 38.38 ± $38\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.26}}$
D-CeLR [dai2024cephalometric] ECCV’24 $5.10\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.16}}$ $27.67\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.28}}$ $60.13\text{\tiny{\color[rgb]{0.55,0.55,0.55}\,$\pm$\,0.27}}$
[…]
Qualitative Evaluation. Fig. 7 shows a visual comparison of our approach with multi-task methods. BlooDet provides greater stability and consistency in detecting bleeding regions and points. In surgical environments with low contrast, competitors tend to be disturbed by surrounding noise. In contrast, our method ensures robust detection across consecutive frames. More visualizations can refer to Supp. G.
5.4 Ablation Analysis
Contributions of Key Component. Table 5 illustrates the contribution of key components in BlooDet. The experimental results show that point memory modeling contributes significantly to detecting bleeding points, e.g., improving PCK-5% by 13.02%. Moreover, the edge generator offers effective edge prompt embedding, enhancing the accuracy of bleeding region detection. In short, each component makes a positive contribution to model performance.
Ablations for Edge Generator. We investigate the effect of edge generator designs in the mask branch. As exhibited in Table 6, the wavelet Laplacian filter shows a strong response to edge clues, mitigating interference from complex backgrounds. Meanwhile, integrating high-resolution features further enhances the quality of edge maps.
Table 5: Ablations for key components of BlooDet for bleeding region and point detection on SurgBlood test set. EG and PMM denote edge generator and point memory modeling modules.
@@ 5.4 Ablation Analysis
Table 6: Influence of edge generator for bleeding region detection.
@@
Table 7: Ablations for optical flow operation via mask maps.
@@
Optical Flow Operation Design. In point memory modeling, we adopt reversed mask maps in conjunction with optical flow maps to estimate the average camera displacement. Thus, we ablate the impact of focusing on different regions in mask maps for bleeding point localization. Table 7 indicates that using the foreground region leads to inferior performance. It may stem from the instability of optical flow in the rapidly changing bleeding areas. In contrast, utilizing the background enables stable motion modeling.
Effect of Mutual Guidance. To validate the effect of our cross-branch mutual guidance, Fig. 9 ablates the impact of point prompt from point branch as well as mask map and mask memory from mask branch. The results indicate that the point prompt contributes to bleeding region detection. For the point branch, the mask maps from previous frames assist with optical flow operation, while mask memory fosters point memory modeling to improve the accuracy of point localization. The attention maps from mask and point decoders visualized in Fig. 8 also verify the effect of mutual guidance. More cross-branch ablations refer to Supp. C.
@@
Figure 9: Effect of mutual guidance between two branches.
@@ 6 Conclusion
This work advances the intelligent bleeding region and point detection in laparoscopic surgery. We contribute an open-source dataset for bleeding detection in real-world surgical videos to facilitate benchmark construction. Accordingly, we design a dual-task synergistic online framework called BlooDet, which assembles mask and point branches in a bidirectional guidance structure, and exploits an edge generator and point memory modeling to enhance the adaptive prompting scheme. Extensive experimental results demonstrate that our method outperforms existing models for both bleeding tasks. We believe this study can facilitate research in intelligent surgical assistance, enhancing intraoperative decision-making and clinical outcomes.